OpenStudy (dj3strella):

i have a question 2. Which is a counterexample that disproves the conjecture? If n is a positive integer, then 2^n – 1 is a prime number. (Points : 1) n = 6 n = 5 n = 3 n = 2 i think the answer is 6

OpenStudy (anonymous):

does that say \[2n-1\] or \[2^{n-1}\]

OpenStudy (dj3strella):

the second one

OpenStudy (anonymous):

then your job is to compute \[2^5-1\\ 2^4-1\\ 2^3-1\\ 2^2-1\] and see which of those is not a prime number

OpenStudy (dj3strella):

its A

OpenStudy (dj3strella):

2^6-1

OpenStudy (anonymous):

\[2^2-1=3\] prime \[2^3-1=7\] prime \[2^4-1=15\] hmmm

OpenStudy (dj3strella):

2^4-1 is not one of them

OpenStudy (anonymous):

oh oops

OpenStudy (anonymous):

\[2^6-1=64-1=63\] yeah, pick that one

OpenStudy (dj3strella):

ok thank you

OpenStudy (anonymous):

A cube has a surface area of one hundred fifty square yards. What is the length of a side? a) five yards b) twelve point five yards c) twenty-five yards d) thirty-seven point five yards

OpenStudy (anonymous):

medal in return

OpenStudy (anonymous):

plzzz

OpenStudy (anonymous):

some one

OpenStudy (dj3strella):

hold on ok

OpenStudy (dj3strella):

@maylineb13

OpenStudy (anonymous):

k

OpenStudy (dj3strella):

the answer is a=5 @maylineb13